OpenStudy (anonymous):

Cylinder A has a radius of 10 inches and a height of 5 inches. Cylinder B has a volume of 750π. What is the percent change in volume between cylinders A and B?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

what's the volume of cylinder A?

OpenStudy (anonymous):

1570.796

OpenStudy (whpalmer4):

keep it symbolic, that will make your life much easier

OpenStudy (anonymous):

what do you mean

OpenStudy (whpalmer4):

instead of multiplying by 3.1415926535..., just write down the symbol for pi

OpenStudy (anonymous):

thats what i did

OpenStudy (whpalmer4):

do the final arithmetic when the problem is done so the volume of cylinder A is 500 pi

OpenStudy (anonymous):

ok so what come next

OpenStudy (whpalmer4):

what is the percentage change between 500 pi and 750 pi? hint: it's the same as the percentage change between 500 and 750, or 50 and 75, or 10 and 15

OpenStudy (whpalmer4):

but if you had 1570.8 and 2356.19, it wouldn't be nearly so obvious that they had pi as a common factor, right?

OpenStudy (anonymous):

these are the answer choices Cylinder B is 50% smaller than cylinder A. Cylinder B is 75% smaller than cylinder A. Cylinder B is 50% bigger than cylinder A. Cylinder B is 200% bigger than cylinder A.

OpenStudy (whpalmer4):

okay, so, cylinder A is 500 pi units, and cylinder B is 750 pi units. is B larger or smaller than A?

OpenStudy (anonymous):

larger

OpenStudy (whpalmer4):

okay, so that eliminates half the answers right there :-) to find the percentage change between A and B, calculate (B - A)/B * 100%

OpenStudy (anonymous):

would it be Cylinder B is 50% bigger than cylinder A.

OpenStudy (whpalmer4):

sorry (B-A)/A * 100% if going from A to B

OpenStudy (whpalmer4):

(750 pi - 500 pi)/(500 pi) * 100% = 250/500 * 100% = 50%, yes.

OpenStudy (anonymous):

ok thanks

OpenStudy (whpalmer4):

when we kept the manipulation symbolic (didn't turn pi into a number), it was much easier and we didn't have to worry about any rounding errors or the like.